Please help me solve this problem... Find the length of the shorter leg of a right triangle if the longer leg is 24 meters and the length of the hypotenuse is 6 meters more than twice the length of the shorter leg.
Answer by cleomenius(959)

24^2 + x^2 = (2x + 6)
576 + x^2 = 4x^2 +24x +36^2
0 = 3x^2 +24x -540 ( we can factor out a 3)
3(x^2 +8x -180) = ( x + 18) ( x - 10)
Therefore, x = 10.
c = 26
24^2 + 10^2 = 26^2
576 m + 100 m = 16^2 = 676 m
Cleomenius.